# From Web Search towards Agentic Deep ReSearch: Incentivizing Search with Reasoning Agents

### v1

@@ 1 Introduction
Recently, test-time scaling (TTS) has emerged as a potent paradigm for boosting the reasoning and agentic capabilities of LLMs (Snell et al., 2024a, ). It assigns additional computation during inference, enabling deeper problem-solving (Zou et al., 2025a, ; Gu et al.,, 2025). Equipped with TTS on reasoning and search, LLMs are set to drive a new search paradigm termed Agentic Deep Research systems, which are capable of autonomous reasoning, on-demand searching, and iterative information synthesis. Demonstrations from deep research products products launched by OpenAI and Google highlight several key advantages of this paradigm: (1) Comprehensive Understanding: Ability to dissect and address complex, multifaceted queries that overwhelm traditional methods (Wei et al.,, 2022); (2) Enhanced Synthesis: Excels at synthesizing information from diverse, potentially conflicting sources into coherent and insightful narratives (Cheng et al.,, 2025); (3) Reduced User Burden: Significantly decreases the cognitive load and manual effort required from users by automating laborious research steps (Sami et al.,, 2024).
Our position is that the LLM-driven Agentic Deep Research framework will inevitably become the dominant paradigm for future information search. In this paper, we comprehensively investigate this transformative shift and make several pivotal contributions: (1) We systematically trace and analyze the evolutionary trajectory of information search paradigms, from traditional keyword-based systems, to conversational LLM chatbots, retrieval-augmented LLMs, and ultimately the Agentic Deep Research; (2) We introduce the test-time scaling (TTS) law for Deep Research, a novel hypothesis formalizing the relationship between allocated inference-time computational resources and the resulting improvements in reasoning and searching tasks; (3) We present extensive evaluations on benchmarks and statistical trends for the open-source implementation to support our position; and (4) As the first to holistically summarize the field of Deep Research, we offer a detailed exploration of critical future research directions, outlining both opportunities and challenges. Our discussion outlines a clear roadmap for this rapidly evolving field, underscoring how this evolution is fundamentally reshaping human interaction with real-world information and guiding subsequent research to further advance and refine Agentic Deep Research.
@@ 3 Towards Agentic Deep Research
The DeepSeek-R1 (Guo et al.,, 2025) represents a significant milestone in this evolution, demonstrating how reinforcement learning can optimize reasoning processes for complex mathematical tasks. By learning from experience and feedback, these systems develop sophisticated strategies for information gathering and synthesis, moving beyond simple pattern matching to true strategic planning. This advancement in reasoning capabilities provides the essential foundation for agentic deep research by enabling systems to autonomously evaluate information needs, strategically decompose complex queries, synthesize information across multiple sources while maintaining logical consistency, and adapt search strategies based on intermediate results and feedback (Jin et al.,, 2025). These capabilities, rooted in advanced reasoning mechanisms, establish the preliminary foundation necessary for conducting deep research tasks that require strategic planning, iterative refinement, and complex decision-making. The integration of reinforcement learning with reasoning frameworks represents a crucial step toward truly agentic systems capable of conducting sophisticated research autonomously, marking a significant advancement from traditional search and retrieval paradigms (Chen et al., 2025a, ).
@@ 3.1 Incentivizing Search with Reasoning Agents
Reinforcement learning (RL) fundamentally changes the search paradigm by letting agents learn through trial and error in interactive environments. Instead of being told how to search, RL-trained agents are incentivized (through feedback or reward functions) to discover, refine, and adapt their reasoning and search strategies for specific goals. Early systems like WebGPT (Nakano et al.,, 2021) and RAG-RL (Huang et al.,, 2025) demonstrated how reward signals (based on human feedback or factual correctness) can guide multi-step retrieval policies that improve response accuracy and trustworthiness. More modular designs like M-RAG (Wang et al., 2024d, ) separate reasoning and retrieval into specialized agents, each trained to collaborate via shared RL objectives. Recent RL-based systems such as Search-R1 (Jin et al.,, 2025), R1-Searcher (Song et al.,, 2025), DeepResearcher (Zheng et al., 2025a, ), ZeroSearch (Sun et al., 2025a, ), and WebAgent-R1 Wei et al., 2025b operate in various search environments from static local corpora and open search APIs to real-world web interfaces. These agents learn to decompose complex tasks, plan query sequences, verify evidence, and adjust their strategies based on environment feedback. Such behaviors are difficult to teach through SFT or prompts alone. Importantly, ReSearch (Chen et al., 2025b, ) and ReARTeR (Sun et al., 2025b, ) go a step further by optimizing not just factual correctness, but also alignment with transparent, interpretable reasoning. ReARTeR introduces a dual-model approach that incentivizes both outcome quality and step-wise explainability, offering a more human-aligned path to trustworthy automation.
@@ 4.3 TTS Trade-offs for Search and Reasoning
Agentic Deep Research systems integrate both search and reasoning, each consuming part of a limited token budget. Under such constraints, a natural trade-off emerges: allocating more tokens to search (e.g., issuing broader or more detailed queries) reduces the capacity available for reasoning (e.g., multi-hop inference or synthesis), and vice versa. This balance is task-dependent. Search-heavy tasks such as multi-hop RAG (Xiong et al.,, 2024; Shao et al.,, 2023) or literature surveys prioritize broader content access, while reasoning-heavy tasks like causal analysis or math verification (Snell et al., 2024b, ) require deeper internal processing. Building systems that adaptively allocate token budgets between search and reasoning based on task characteristics is critical for maximizing effectiveness and efficiency in Agentic Deep Research. We anticipate the emergence of a test-time scaling law that governs the optimal balance between search and reasoning under different task conditions. Furthermore, training models that can dynamically allocate and manage budget across these two components is a defining capability of next-generation deep research systems.

### v2

@@ 1 Introduction
Recently, test-time scaling (TTS) has emerged as a potent paradigm for boosting the reasoning and agentic capabilities of LLMs (Snell et al.,, 2024). It assigns additional computation during inference, enabling deeper problem-solving (Zou et al., 2025a, ; Gu et al.,, 2025). Equipped with TTS on reasoning and search, LLMs are set to drive a new search paradigm termed Agentic Deep Research systems, which are capable of autonomous reasoning, on-demand searching, and iterative information synthesis. Demonstrations from deep research products launched by OpenAI and Google highlight several key advantages of this paradigm: (1) Comprehensive Understanding: Ability to dissect and address complex, multifaceted queries that overwhelm traditional methods (Wei et al.,, 2022); (2) Enhanced Synthesis: Excels at synthesizing information from diverse, potentially conflicting sources into coherent and insightful narratives (Cheng et al.,, 2025); (3) Reduced User Burden: Significantly decreases the cognitive load and manual effort required from users by automating laborious search steps (Sami et al.,, 2024).
Our position is that the LLM-driven Agentic Deep Research framework will inevitably become the dominant paradigm for future information search. In this paper, we comprehensively investigate this paradigm shift and make four key contributions: (1) We systematically trace and analyze the evolutionary trajectory of information search paradigms, from traditional keyword-based search, through conversational LLM chatbots and naive search-augmented LLMs, and ultimately to the Agentic Deep Research; (2) We introduce the test-time scaling (TTS) law for Deep Research, a novel hypothesis formalizing the relationship between inference-time computational resources allocation and the resulting improvements in LLMs reasoning capabilities and knowledge depth; (3) We conduct extensive evaluations on existing Agentic Deep Research models, coupled with analyses of open-source implementations to support our position; and (4) As the first to holistically summarize the field of Deep Research, we offer a detailed exploration of critical future research directions, outlining both opportunities and challenges. Our discussion outlines a clear roadmap for this rapidly evolving field, underscoring how this evolution is fundamentally reshaping human interaction with real-world information and guiding further advancement on Agentic Deep Research.
@@ 3 Towards Agentic Deep Research
The DeepSeek-R1 (Guo et al.,, 2025) represents a significant milestone in this evolution, demonstrating how reinforcement learning can optimize reasoning processes for complex mathematical tasks. By learning from experience and feedback, these systems develop sophisticated strategies for information gathering and synthesis, moving beyond simple pattern matching to true strategic planning. This advancement in reasoning capabilities provides the essential foundation for agentic deep research by enabling systems to autonomously evaluate information needs, strategically decompose complex queries, synthesize information across multiple sources while maintaining logical consistency, and adapt search strategies based on intermediate results and feedback (Jin et al.,, 2025). These capabilities, rooted in advanced reasoning mechanisms, establish the preliminary foundation necessary for conducting deep research tasks that require strategic planning, iterative refinement, and complex decision-making. The integration of reinforcement learning with reasoning frameworks represents a crucial step toward truly agentic systems capable of conducting sophisticated research autonomously, marking a significant advancement from traditional search and retrieval paradigms (Chen et al., 2025b, ).
@@ 3.1 Incentivizing Search with Reasoning Agents
Reinforcement learning (RL) fundamentally changes the search paradigm by letting agents learn through trial and error in interactive environments. Instead of being told how to search, RL-trained agents are incentivized (through feedback or reward functions) to discover, refine, and adapt their reasoning and search strategies for specific goals. Early systems like WebGPT (Nakano et al.,, 2021) and RAG-RL (Huang et al., 2025a, ) demonstrated how reward signals (based on human feedback or factual correctness) can guide multi-step retrieval policies that improve response accuracy and trustworthiness. More modular designs like M-RAG (Wang et al., 2024d, ) separate reasoning and retrieval into specialized agents, each trained to collaborate via shared RL objectives. Recent RL-based systems such as Search-R1 (Jin et al.,, 2025), R1-Searcher (Song et al.,, 2025), DeepResearcher (Zheng et al., 2025a, ), ZeroSearch (Sun et al., 2025a, ), and WebAgent-R1 Wei et al., 2025b operate in various search environments from static local corpora and open search APIs to real-world web interfaces. These agents learn to decompose complex tasks, plan query sequences, verify evidence, and adjust their strategies based on environment feedback. Such behaviors are difficult to teach through SFT or prompts alone. Importantly, ReSearch (Chen et al., 2025b, ) and ReARTeR (Sun et al., 2025b, ) go a step further by optimizing not just factual correctness, but also alignment with transparent, interpretable reasoning. ReARTeR introduces a dual-model approach that incentivizes both outcome quality and step-wise explainability, offering a more human-aligned path to trustworthy automation.
@@ 4.3 TTS Trade-offs for Search and Reasoning
Agentic Deep Research systems integrate both search and reasoning, each consuming part of a limited token budget. Under such constraints, a natural trade-off emerges: allocating more tokens to search (e.g., issuing broader or more detailed queries) reduces the capacity available for reasoning (e.g., multi-hop inference or synthesis), and vice versa. This balance is task-dependent. Search-heavy tasks such as multi-hop RAG (Xiong et al.,, 2024; Shao et al.,, 2023) or literature surveys prioritize broader content access, while reasoning-heavy tasks like causal analysis or math verification (Snell et al.,, 2024) require deeper internal processing. Building systems that adaptively allocate token budgets between search and reasoning based on task characteristics is critical for maximizing effectiveness and efficiency in Agentic Deep Research. We anticipate the emergence of a test-time scaling law that governs the optimal balance between search and reasoning under different task conditions. Furthermore, training models that can dynamically allocate and manage budget across these two components is a defining capability of next-generation deep research systems.
@@ Appendix B Open-Source Deep Research Implementations
IKEA (Huang et al., 2025b, ) Qwen RL NQ, HotpotQA NQ, HotpotQA, PopQA, 2Wikimultihopqa Link